# Over-parameterization and Adversarial Robustness in Neural Networks: An Overview and Empirical Analysis

### v1

@@ 3.4 How reliable are these evaluations?
In an attempt to verify the generalizability of the models, different works carried out analysis under different scenarios, ex., real-world environment, adversarial environment, noisy environment, etc. For safety- and security-sensitive tasks, robustness under an adversarial environment is of utmost importance.
Various earlier works  [38, 39, 31] have shown that constructing effective adversarial defense is harder than expected. Many works that propose new defenses, to evaluate the system security, use suboptimal attack methods, often leading to a false sense of securities. Even the use of simple gradient-based attacks like PGD, as observed in  [40, 41] can fail under certain parameter choices leading to over-estimation of robustness. As explained in the previous section, the related works (e.g., [32, 26]), have employed PGD to compare the robustness of neworks with different sizes, but they have not verified if the attack was succeding. Therefore, it is necessary to verify the reliability of the security evaluations. In this work, we verified our results using Indicators of Attack Failures (IoAF) presented by Pintor et. al  [31]. We also used AutoAttack as an additional ensemble of attacks to measure the reliability of the results shown in section 4.
@@
With over decade-long research on devising new attacks and failures to study the robustness of models, many recent works have shown that the attacks may fail to generate viable adversarial samples, for example, depending on the choice of loss function $L$ and the target model $\theta$, some loss functions may exhibit obfuscated gradients [42, 38] which prevents gradient-based attacks from generating an optimal attack even when they exist in the feasible domain. Another known issue in constructing attacks is attack-optimization failures which are observed at the time of running gradient-based attacks  [11, 43], which may be either due to usage of suboptimal parameters or inappropriate implementation. Therefore we use IoAF [31] which is an automated testing package aimed at detecting known failures within the optimization process of gradient-based attacks. We use IoAF to make sure the attacks generated on over-parameterized models are reliable and the final empirical observations are unbiased.
Most models evaluations are relied on the results of PGD or its weaker veriosn: FGSM [5]. Models evaluated on single or weak attacks can run into the problem of suboptimal solution found on a local minima that maybe far from the global optimum. Therefore, for reliable evaluation of the over-parameterized models, we run a diverse attack to increase the probability to find better approximation of the global optimim. We use AutoAttack to add diversity in the evaluation. Their parameter-free, computationally cheap and user-independent parameters make it a useful tool for measuring the reliability of the security evaluations of the model (explained in detail in subsubsection 3.2.2).
@@ 4.1 Experimental setup
In this work, we consider two datasets that have been widely used in the pre-existing work in this field. The MNIST dataset  [44] is a handwritten grayscale digit from 0 to 9. It consists of 70,000 28x28 images. The CIFAR-10 dataset  [45] is a colored image of ten different objects. It consists of 60,000 32x32 images.
We used fixed number of training and test samples throughout all the models; MNIST: 3000 training samples and 1000 test samples, CIFAR10: 20,000 training samples and 1000 test samples. The goal of the experiments is to study the impact of over-parameterization and their security, therefore, we keep all parameters fixed and vary the model size.
The three considered models are: i) Convolutional Neural Network (CNN), ii) a fully connected model with Rectified Linear activation (FC-ReLU), iii Residual Network (ResNet)  [37] The first two models are trained on MNIST and ResNet is trained on CIFAR10. Having variety in models make sure that the trend is not dependent on model type but on model size.
We first train the models with training data, evaluate them with benign test samples, and refer to them as base performance. These evaluations are marked as $\epsilon=0.0$ in the plots. We then construct adversarial samples using PGD confined to specified $l_{2}$-ball on these models. Attacks are generated with varying degrees of perturbations, for MNIST-CNN: $\epsilon=\{0.3,0.98,1.65,2.33,3.0\}$, MNIST-FCRELU: $\epsilon=\{0.1,0.45,0.80,1.15,1.5\}$ and for CIFAR-10: $\epsilon=\{0.01,0.08,0.15,0.23,0.30\}$. We use 100 iterations and a step-size of 1e-3 for both MNIST and CIFAR10.
We leverage SecML and Pytorch libraries to implement the attacks. Further, to employ the AutoAttack framework, we leveraged the AutoAttack library by Croce et al.  [33]
@@ 4.2.1 Base Performance
From Figure 2  we see the base performance, namely the performances on the clean (unperturbed) test data, of the three configurations considered in this work: CNN trained on MNIST, FC-ReLU trained MNIST, and Resnet18 trained CIFAR10 respectively. We can observe that the classifier’s performance changes when we increase the number of network parameters. None of the performance on the test data worsens as would have been expected from the classical theory of overfitting for largely parameterized models. Instead, the error rates continue to decrease for over-parameterized networks supporting the generalizability claims. This is a well-known result, but it serves to us as baseline. It’s worth noting that the focus of the work is not on acquiring the best possible classification performance on the base model but instead how the robustness of the model varies as we increase the size of the model.
@@ 4.2.2 Model Robustness
The Figure 3 shows the security evaluation curves for networks attacked with PGD attack with $l_{2}$-norm constraint on an increasing number of parameters and various degrees of perturbations. While it is understood that the error rate would increase with increasing degrees of perturbations, it is interesting to note that error rates for the over-parameterized model are significantly lower than that of the models with lesser parameters. We see similar trends for all dataset-architecture pairs. This is an important observation as it suggests that over-parameterized models are not only generalizable but are also robust against adversarial samples. However, this outcome may be due to a failure of the PGD attack used for the evaluation.
@@ 4.2.3 Reliability of Model Robustness
The results in Figure 4 include all dataset-architecture configurations tested employing AutoAttack. The trends follow that of PGD from subsubsection 4.2.2. From these performance plots, we can say that even though one could train a close-to-perfect model on the unperturbed test set using smaller networks since there isn’t a significant decline in the benign data performance, over-parameterized models significantly help with the robustness. We also performed IoAF and observed no triggers for attack failures for any of the six indicators. These results provide reliability to the observed higher robustness of overparametrized networks.
@@ 5 Conclusions and Future Works
In this work, we empirically study how robust over-parameterized models are in adversarial environments as compared to their under-parameterized counterparts. These impressive generalizable properties of large models defy the widely accepted understanding of the model selection method that uses the Empirical Risk Minimisation (ERM) principle that is biased to select the smallest network that can fit the data well, is often on the left side of the second descent of the double descent curve. From the classical statistical notion of bias-variance tradeoff, the optimal model is neither too large nor too small to avoid overfitting and underfitting, respectively. Yet the performance of large models outperforms that of smaller models, making their usage ubiquitous. However, when it comes to the security of the large models, the literature has contradictory claims in favor and against the robustness in an adversarial environment. Most of the literature in the domain does not verify the reliability of the attacks either.
Our results show that over-parameterized models not only generalize well but are also more robust to carefully crafted adversaries in their environment than their under-parametrized counterparts. We provide the empirical evidence to show that even when increasing the number of parameters is no more beneficial for the accuracy on clean data, it is beneficial for the robustness against adversarial examples.
Future directions of this work may include evaluating over-parameterized models against privacy attacks. It would also be interesting to investigate if it is possible to obtain smaller networks with accuracy and robustness comparable to those of over-parameterized networks to avoid computational burden.
6 Acknowledgement
@@
This work has been carried out while Srishti Gupta was enrolled in the Italian National Doctorate on Artificial Intelligence run by Sapienza University of Rome in collaboration with University of Cagliari.

### v2

@@ 3.4 How reliable are these evaluations?
In an attempt to verify the generalizability of the models, different works carried out analysis under different scenarios, ex., real-world environment, adversarial environment, noisy environment, etc. For safety- and security-sensitive tasks, robustness under an adversarial environment is of utmost importance. Various earlier works  [38, 41, 33] have shown that constructing effective adversarial defense is harder than expected. Many works that propose new defenses, to evaluate the system security, use suboptimal attack methods, often leading to a false sense of securities. Even the use of simple gradient-based attacks like PGD, as observed in  [42, 43] can fail under certain parameter choices leading to over-estimation of robustness. As explained in the previous section, the related works (e.g., [34, 28]), have employed PGD to compare the robustness of neworks with different sizes, but they have not verified if the attack was succeding. Therefore, it is necessary to verify the reliability of the security evaluations. In this work, we verified the results obtained with the PGD attack using Indicators of Attack Failures (IoAF) presented by Pintor et. al  [33]. We also used AutoAttack as an additional ensemble of attacks to measure the reliability of the results shown in Sect. 4. With over decade-long research on devising new attacks and failures to study the robustness of models, many recent works have shown that the attacks may fail to generate viable adversarial samples, for example, depending on the choice of loss function $L$ and the target model $\theta$, some loss functions may exhibit obfuscated gradients [44, 38] which prevents gradient-based attacks from generating an optimal attack even when they exist in the feasible domain. Another known issue in constructing attacks is attack-optimization failures which are observed at the time of running gradient-based attacks  [11, 45], which may be either due to usage of suboptimal parameters or inappropriate implementation. Therefore we use IoAF [33] which is an automated testing package aimed at detecting known failures within the optimization process of gradient-based attacks. We use IoAF to make sure the attacks generated on over-parameterized models are reliable and the final empirical observations are unbiased. Most models evaluations are relied on the results of PGD or its weaker veriosn: FGSM [5]. Models evaluated on single or weak attacks can run into the problem of suboptimal solution found on a local minima that maybe far from the global optimum. Therefore, for reliable evaluation of the over-parameterized models, we run a diverse attack to increase the probability to find better approximation of the global optimim. We use AutoAttack to add diversity in the evaluation. Their parameter-free, computationally cheap and user-independent parameters make it a useful tool for measuring the reliability of the security evaluations of the model (explained in detail in Sect. 3.2.2).
Figure 2: Base performance of trained classifiers against the parameter size: (a) CNN trained on MNIST, (b) linear classifier trained on MNIST, and (c) ResNet model trained on CIFAR10.
@@ 4.1 Experimental setup
In this work, we consider two datasets that have been widely used in the pre-existing work in this field. The MNIST dataset  [46] is a handwritten grayscale digit from 0 to 9. It consists of 70,000 28x28 images. The CIFAR10 dataset  [47] is a colored image of ten different objects. It consists of 60,000 32x32 images. We used fixed number of random training and test samples throughout all the models; MNIST: 3000 training samples and 1000 test samples, CIFAR10: 20,000 training samples and 1000 test samples. These samples are then loaded sequentially in the train and test loaders. The goal of the experiments is to study the impact of over-parameterization and their security, therefore, we keep all parameters fixed and vary the model size.
Let $\mathbf{x}\in\mathcal{X}$ and $y\in\mathcal{Y}$ be the input and output space respectively and $\mathcal{P}$ be a distribution over $\mathcal{X}\times\mathcal{Y}$. The models have input dimension d, i.e. $\mathbf{x}\in\mathbb{R}^{d}$, output dimension c, i.e. $y\in\mathbb{R}^{c}$, and number of hidden units m. For example, for a 2 layered network, the output of the network would be $f_{h^{{}^{\prime}},h^{{}^{\prime\prime}}}(\mathbf{x})=h^{{}^{\prime\prime}}(h^{{}^{\prime}}(\mathbf{x}))$, such that $h^{{}^{\prime}}\in\mathbb{R}^{m\times d}$ and $h^{{}^{\prime\prime}}\in\mathbb{R}^{c\times m}$ where $h^{{}^{\prime}},h^{{}^{\prime\prime}}$ are hidden layers. Using the given setting, we considered the following three models taking inspiration from [40]:
  • CNN: This Convolutional Neural Network has 2 convolutional layers, the first having a number of neurons equal to z and the second having 2 * z neurons. We trained different architectures, setting z equal to [1, 2, 4, 6, 8, 10, 15, 20, 25, 30] obtaining a total of 10 architectures with parameter size varying from 760 to 38170.
  • FC-ReLU: A fully connected network with rectified linear activation that has two hidden layers, the first having a number of neurons equal to z and the second having 8 * z neurons. We trained different architectures setting z equal to [4, 6, 8, 10, 15, 20, 25, 30, 35, 40], obtaining a total of 10 architectures with parameter sizes varying from 3630 to 47730.
  • ResNet: The Residual Network consists of a convolution layer followed by 6 residual blocks. We trained different architectures, setting number of neurons z equal to [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] obtaining a total of 10 architectures with parameter size varying from 25k to 11 million expanding number of channels, say k, in the first convolution layer such that k = ${2}^{2+z/2}]$.
The first two models (CNN and FC-ReLU) are trained on MNIST whereas, ResNet is trained on CIFAR10. We trained by minimizing the cross entropy loss. We use the Adam optimizer with a learning rate of 0.001 and batch size of 20 and 50 for MNIST and CIFAR10 training respectively. Having variety in models make sure that the trend is not dependent on model type but on model size. We first train the models with training data, evaluate them with benign test samples, and refer to them as base performance. These evaluations are marked as $\epsilon=0.0$ in the plots. We then construct adversarial samples using PGD confined to specified $l_{2}$-ball on these models. Attacks are generated with varying degrees of perturbations, for MNIST-CNN: $\epsilon=\{0.3,0.98,1.65,2.33,3.0\}$, MNIST-FC-ReLU: $\epsilon=\{0.1,0.45,0.80,1.15,1.5\}$ and for CIFAR10: $\epsilon=\{0.01,0.08,0.15,0.23,0.30\}$. We use 100 iterations and a step-size of 1e-3 for both MNIST and CIFAR10. We leverage SecML and Pytorch libraries to implement the attacks. Further, to employ the AutoAttack framework, we leveraged the AutoAttack library by Croce et al.  [35]
@@ 4.2.1 Base Performance
From Fig. 2  we see the base performance, namely the performances on the clean (unperturbed) test data, of the three configurations considered in this work: CNN trained on MNIST, FC-ReLU trained MNIST, and Resnet18 trained CIFAR10 respectively. We can observe that:
  • The classifier’s performance changes when we increase the number of network parameters;
  • None of the performance on the test data worsens as would have been expected from the classical theory of overfitting for largely parameterized models. Instead, the error rates continue to decrease for over-parameterized networks supporting the generalizability claims.
This are well-known results, but they serves to us as baseline. It’s worth noting that the focus of the work is not on acquiring the best possible classification performance on the base model but instead how the robustness of the model varies as we increase the size of the model.
@@ 4.2.2 Model Robustness
The Fig. 3 shows the security evaluation curves (SEC) for networks attacked with PGD attack with $l_{2}$-norm constraint on an increasing number of parameters and various degrees of perturbations. While it is understood that the error rate would increase with increasing degrees of perturbations, it is interesting to note that error rates for the over-parameterized model are significantly lower than that of the models with lesser parameters. We see similar trends for all dataset-architecture pairs. This is an important observation as it suggests that over-parameterized models are not only generalizable but are also robust against adversarial samples. However, this outcome may be due to a failure of the PGD attack used for the evaluation.
Figure 3: SEC plots of over-parameterized models using PGD $l_{2}$ attack on (a) CNN trained on MNIST, (b) FC-ReLU trained on MNIST, and (c) ResNet trained on CIFAR10. Each curve represents a network trained with the number of parameters specified in the legend.
Figure 4: SEC plots of over-parameterized models using AutoAttack on (a) CNN trained on MNIST, (b) FC-ReLU trained on MNIST, and (c) ResNet trained on CIFAR10. Each curve represents a network trained with the number of parameters specified in the legend.
@@ 4.2.3 Reliability of Model Robustness
The results in Fig. 4 include all dataset-architecture configurations tested employing AutoAttack. Results:
  • Even though one could train a close-to-perfect model on the unperturbed test set using smaller networks since there isn’t a significant decline in the benign data performance, over-parameterized models significantly help with the robustness.
  • The trends obtained with AutoAttack follow that of PGD from Sect. 4.2.2; moreover we have verified the affidability of the Sect. 4.2.2 plots performing IoAF and we have observed no triggers for attack failures for any of the six indicators. These results verify that the analysis performed in this paper is well-conducted, providing reliability to the observed higher robustness of overparameterized networks.
@@ 5 Conclusions and Future Works
In this work, we empirically study how robust over-parameterized models are in adversarial environments as compared to their under-parameterized counterparts. These impressive generalizable properties of large models defy the widely accepted understanding of the model selection method that uses the Empirical Risk Minimization (ERM) principle that is biased to select the smallest network that can fit the data well, is often on the left side of the second descent of the double descent curve. From the classical statistical notion of bias-variance tradeoff, the optimal model is neither too large nor too small to avoid overfitting and underfitting, respectively. Yet the performance of large models outperforms that of smaller models, making their usage ubiquitous. However, when it comes to the security of the large models, the literature has contradictory claims in favor and against the robustness in an adversarial environment. This claims are made without verifying the reliability of the attacks and, for this reason, they are unreliable. Our results conducted with an attack of which we have verified the effectiveness, show that over-parameterized models not only generalize well but are also more robust to carefully crafted adversaries in their environment than their under-parameterized counterparts. We provide the empirical evidence to show that even when increasing the number of parameters is no more beneficial for the accuracy on clean data, it is beneficial for the robustness against adversarial examples. Future directions of this work may include evaluating over-parameterized models against privacy attacks. It would also be interesting to investigate if it is possible to obtain smaller networks with accuracy and robustness comparable to those of over-parameterized networks to avoid computational burden.
6 Expected Impact on Future Work
While some work about the robustness of over-parameterized deep neural networks was present in the literature, their claim was contradictory, and their analysis was unreliable because the effectiveness of the performed attacks was not verified. The analysis presented in this work is verified with tools ad-hoc to analyze its reliability and shows that over-parameterized networks are more robust. We believe the empirical evidence provided by this manuscript is the starting point for theoretical works explaining why over-parametrized networks are more robust. Moreover, we believe that adding a spotlight on the reliability of the robustness analysis of over-parameterized networks while highlighting the incoherence of the previously proposed analysis will lead to more reliable future works on this topic.
7 Funding
@@
This work has been carried out while Srishti Gupta was enrolled in the Italian National Doctorate on Artificial Intelligence run by Sapienza University of Rome in collaboration with the University of Cagliari. This work was also partially supported by project FAIR (PE00000013) under the NRRP MUR program funded by the EU – NGEU (CUP: J23C24000090007).
8 Author contributions
S.G. wrote the manuscript draft, performed the experiments, and made some updates to the code written by L.D. Z.C. made the initial exploratory experiments. L.D. wrote the code for the experiments currently reported in the paper. A.D. supervised Z.C. and S.G. B.B., X.F., and F.R. acquired funding. All authors reviewed the manuscript.
9 Declarations
9.1 Data availability
This manuscript does not report data generation and utilizes publicly available datasets.
9.2 Ethic Declaration
Not Applicable
9.3 Consent to Participate
Not Applicable
9.4 Clinical Trial Number
Not Applicable
9.5 Consent to Publish
Not Applicable
9.6 Competing Interest
The authors declare no competing interests.
9.7 Declaration of Generative AI in Scientific Writing
During the preparation of this work, the authors used Grammarly in order to edit and refine the language. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.